# Algorithmic resolution of crowd-sourced moderation on X in polarized settings across countries

### v1

@@ G.1.1 United States
Specifically, we consider the 109 US news media outlets, used in at least 100 Community Notes, whose ideological skew was also evaluated by Ad Fontes Media  (?). We subsequently compare the Left-Right bias of the news outlets, as assessed by Ad Fontes Media, and the average latent ideology $\theta_{r}$ of the Community Notes contributors who authored notes using these media as sources. As displayed in Figure S29, we observe a strong Pearson correlation between these quantities: $\rho=0.744$ $p<10^{-4}$, with a $[0.65,~{}0.82]$ Confidence Interval at 95% for correlation, estimated via Fisher transformation.
@@
Notes incorporating expert fact-checking sources reach Helpful Status at significantly higher rates ($21.3\%$
$[18.3\%,~{}24.1\%]$) than average, with this effect more pronounced for notes authored by Right-leaning contributors ($\theta_{r}>0$): $26.9\%[23.3\%,~{}30.8\%]$ versus $17.6\%~{}[14.0\%,~{}20.7\%]$ for Left-leaning contributors ($\theta_{r}<0$); 95% confidence interval over 100 bootstraps with replacement over the notes and the set of fact-checking organizations considered in the analysis.
@@ G.2 Anti-Elite dimension
Figure S31: X accounts positioned in the (Left-Right; Anti-Elite) 2D plane, segmented by the majority sign of the latent ideology $\theta_{n}$ of their associated Community Notes (red for $\theta_{n}<0$ and blue for $\theta_{n}>0$). The direction that best separates notes by majority outcome, learned through logistic regression, is displayed as $CN$ along with the corresponding AUC score for sign prediction based on account positions in the 2D plane (standard deviation from 10-fold cross-validation). The direction structuring each country’s political landscape $\delta_{1}$ is projected onto the plane.

### v2

@@ G.1.1 United States
Specifically, we consider the 109 US news media outlets, used in at least 100 Community Notes, whose ideological skew was also evaluated by Ad Fontes Media (?). We subsequently compare the Left-Right bias of the news outlets, as assessed by Ad Fontes Media, and the average latent ideology $\theta_{r}$ of the Community Notes contributors who authored notes using these media as sources. As displayed in Figure S29, we observe a strong Pearson correlation between these quantities: $\rho=0.744$ $p<10^{-4}$, with a [0.65,0.82][0.65,~0.82] Confidence Interval at 95% for correlation, estimated via Fisher transformation.
@@
Notes incorporating expert fact-checking sources reach Helpful Status at significantly higher rates ($21.3\%~[18.3\%,~24.1\%]$) than average, with this effect more pronounced for notes authored by Right-leaning contributors ($\theta_{r}>0$): $26.9\%~[23.3\%,~30.8\%]$ versus $17.6\%~[14.0\%,~20.7\%]$ for Left-leaning contributors ($\theta_{r}<0$); 95% confidence interval over 100 bootstraps with replacement over the notes and the set of fact-checking organizations considered in the analysis.
@@ G.2 Anti-Elite dimension
Figure S31: X accounts positioned in the (Left-Right; Anti-Elite) 2D plane, segmented by the majority sign of the latent ideology $\theta_{n}$ of their associated Community Notes (red for $\theta_{n}<0$ and blue for $\theta_{n}>0$). The direction that best separates notes by majority outcome, learned through logistic regression, is displayed as CNCN along with the corresponding AUC score for sign prediction based on account positions in the 2D plane (standard deviation from 10-fold cross-validation). The direction structuring each country’s political landscape $\delta_{1}$ is projected onto the plane.